Charleston Company has elected to use the dollar-value LIFO retail method to value its inventory. The following data has been accumulated from the accounting records:

Cost Retail
Merchandise inventory, January 1, 2018 $ 333,840 $ 520,000
Net purchases 674,928 1,022,000
Net markups 14,200
Net markdowns 4,200
Net sales 661,000
Pertinent retail price indexes:
January 1, 2018 1.00
December 31, 2018 1.10


Required:
Estimate the ending inventory for December 31, 2018. (Round "Cost-to-retail percentage" to two decimal place.)

Estimated ending inventory (At Retail)?

Estimated ending inventory (At Cost)?
Answer #1
Computation of Estimated the ending inventory for December 31, 2018
Cost ($) Retail ($)
Inventory Jan. 1     333,840.00     520,000.00
Net purchases     674,928.00 1,022,000.00
Net markups       14,200.00
Net markdowns       (4,200.00)
Goods available for sale (excluding beginning inventory)     674,928.00 1,032,000.00
Goods available for sale (including beginning inventory) 1,008,768.00 1,552,000.00
Net sales    (661,000.00)
Ending inventory, at retail     891,000.00
Cost-to-retail ratio = 674.928/1,032,000 = 65%
Estimated ending inventory (At Retail) = 891,000 / 1.10 = $ 810,000
Estimated ending inventory (At Cost)
520,000 x 1.00 = $520,000
290,000 x 1.10 = $319,000
Total Estimated ending inventory (At Cost) = $839,000